# IDEA: An Inverse Domain Expert Adaptation Based Active DNN IP Protection Method

### v1

@@ 8       $\overline{D}_{\text{tr},j}=D_{\text{tr}}+D^{+}_{\text{tr},j}+D^{-}_{\text{tr},j}$
10       Fine-tune the real expert model for $E_{1}$ epochs from the original unprotected model. $\triangleright$ Eq. (17)
11       Separately optimize two randomly initialized models for $E_{2}$ epochs to obtain the fake experts. $\triangleright$ Eq. (18) (19)
@@
13      Construct the MoE model. $\triangleright$ Eq. (20)
14
@@
16       Distill the MoE teacher model for $E_{3}$ epochs to generate the protected student model $f_{\theta^{p}_{j}}(\cdot)$. $\triangleright$ Eq. (21)
17
18      return $f_{\theta^{p}_{j}}(\cdot)$.
Distribute each model-key pair and the encoder (i.e., $\{f_{\theta^{p}_{j}}(\cdot),k_{j}\}$ and $\mathcal{E}(\cdot,\cdot)$) to the corresponding legal user.
@@
IV-D Ownership Verification and Culprit Tracing
Intercepted image decoding. An adversary can be tracked by extracting the key from queries submitted to the suspected model by using the SteganoGAN decoder, which is kept private by the model owner. If the calculated Hamming distance (HD) between the key $\hat{s}$ extracted from the intercepted test image and each user key recorded by the model owner exceeds a predetermined threshold, the user of that recorded key is identified as the culprit.
Given a set of intercepted images $D_{\text{int}}$ and a threshold $\epsilon_{3}$, we define the tracing success rate (TSR) as
$\text{TSR}(k,D_{\text{int}})=\mathbb{E}_{x\sim D_{\text{int}}}\left[\mathbb{I}\left[\text{HD}\left(k,\text{Vote}\left(\text{Split}(\mathcal{D}(x))\right)\leq\epsilon_{3}\right)\right]\right],$ (27)
where $k$ denotes a recorded user key. $\mathbb{I}\left[\cdot,\cdot\right]$ is another indicator function which returns 1 if the HD between the recorded and decoded keys does not exceed $\epsilon_{3}$, and 0 otherwise.
Unfortunately, this tracing method requires the owner to intercept key-embedded input samples before the inference stage, which can be difficult sometimes. Moreover, the ownership of the suspected model cannot be attested.
Black-box query. The owner can also extract the user key from a black-box suspected model by performing a series of queries and analyzing the outputs. Specifically, an innocent model learns nothing about the user key will maintain the same high accuracy on both benign and key-embedded test samples. A pirated or redistributed model performs well only on authorized images encoded with a valid key. The dishonest user can be identified by the correct key that unlocks the original performance of the misappropriated model.
@@ V-A1 Datasets and Networks
Table I specifies the five public image classification datasets, CIFAR-10 [42], CIFAR-100 [42], GTSRB [43], Caltech-101 [44], and ImageNet [45], with different resolutions and numbers of classes. To speed up the training without loss of generality, we perform the ImageNet evaluation on its subset of 10 arbitrarily selected classes named ImageNet-10. The real and fake experts share the same architecture across ResNet-18 (Res-18) [46], PreActResNet-18 (PreRes-18) [46], SENet [47] and MobileNet-V2 (Mob-V2) [48].
TABLE I: Specifications of the experimental datasets.
Dataset # Labels Input Size # Training / Test Images
CIFAR-10 10 32 $\times$ 32 50000 / 10000
CIFAR-100 100 32 $\times$ 32 50000 / 10000
GTSRB 43 32 $\times$ 32 39208 / 12630
Caltech-101 101 32 $\times$ 32 6942 / 1735
ImageNet-10 10 224 $\times$ 224 12000 / 1000
V-A2 Key size
We set $r$ according to the input size of the dataset. For ImageNet-10, $r$ is set to 32, which gives a key space of $2^{32\times 32}$. The $32\times 32$ key is replicated 49 times and then squarely concatenated to the same size of $224\times 224$ as the benign images before they are input to the SteganoGAN encoder. For the other four datasets, we set $r$ to 16 with a key space of $2^{16\times 16}$. The $16\times 16$ key is replicated four times and concatenated to the size of $32\times 32$.
V-A3 Other settings
For each classification task, a dedicated SteganoGAN model is trained using 2500 samples randomly selected from the benign training dataset. The unprotected model $f_{\theta}(\cdot)$ is trained from scratch for 200 epochs, and from which the real expert $f_{\theta^{R}_{j}}(\cdot)$ is fine-tuned for 40 epochs, both using the Stochastic Gradient Descent optimizer. Each fake expert is separately trained for 1000 iterations. During the knowledge distillation process, the student model is optimized from scratch for 50 epochs using the Adam optimizer. Ten protected models are generated from the same source model with different random keys for distribution to ten different users. Several common data augmentations, including random cropping, random horizontal flipping, random rotation and random erasing, are applied during training to alleviate overfitting.
@@ V-A3 Other settings
V-B Evaluation
V-B1 Stealthiness
Fig. 4 presents some examples of benign and authorized images for CIFAR-10, Caltech-101, GTSRB, and ImageNet-10 datasets. We also provide the magnified (×$\times 10$) residual images obtained by the absolute difference in pixel values between the benign and authorized images. The perturbations introduced by SteganoGAN encoding are visually imperceivable. In other words, the attacker can hardly figure out the embedded key by inspecting a few test images encoded with the same key. Furthermore, the perturbations are sample-specific. Even if the attacker can compute a residual image from a pair of benign and authorized image of a protected model, it is impossible to generate an authorized image by adding the residual image to a different test image. The means and standard deviations of three objective image quality assessments (IQA), Structural Similarity Index Measure (SSIM) [49], Peak Signal to Noise Ratio (PSNR) and Learned Perceptual Image Patch Similarity (LPIPS) [50] are presented in Table II. The outstanding IQA scores prove the stealthiness of the key-encoded images.
@@ V-B2 Effectiveness and fidelity
The inference performances of both unprotected and protected models on benign, authorized, and noise test images are presented in Table III. The results indicate that unprotected models consistently achieve high prediction accuracy across all three domains. All protected models fulfill the fidelity requirement of Eq. (4), with inference accuracy of authorized images comparable to the benign accuracy of the corresponding unprotected model. The accuracy drop is no more than 1.5%. Additionally, the protected models perform poorly on benign and noise images, with near-random guessing accuracies, making them useless to unauthorized users without the correct key.
@@ V-B3 Uniqueness
(a) CIFAR-10
(b) CIFAR-100
(c) Caltech-101
(d) GTSRB
(e) ImageNet-10
Figure 5: Uniqueness evaluation of IDEA. Row 1: confusion matrices for the evaluation of 10 different keys on the 10 encoded models. ResNet-18 is the architecture used for all the protected models. Row 2: Accuracies of the protected models on test images embedded with incorrect keys with a few erroneous bits.
We trained 10 protected models encoded with different keys and evaluated the prediction performance across all possible combinations of model-key pairs. The confusion matrices are presented in the first row of Fig. 5, where each row denotes a protected model and each column denotes a key. The diagonal and off-diagonal values represent test accuracies for matched and mismatched model-key pairs, respectively. All but the diagonal values of these confusion matrices are very low (<20%), indicating that user keys among protected models are non-exchangeable. Therefore, the protected model of one legitimate user cannot achieve accurate prediction on encoded images generated with another legitimate user’s key.
An attacker may have stolen a protected model and the SteganoGAN encoder from a legitimate user. Without the legitimate user key, the complexity of brute force key guessing is $O(2^{r^{2}})$, provided that the protected model produces random output to any input images generated by the stolen encoder with the wrong guessed key even if the guessed key contains only one error bit. To evaluate the tolerance of the protected model to bit errors in the key used for encoding the test images, we randomly flipped a few bits, ranging from 1 to 24, of the user keys. As shown in the second row of Fig. 5, for the four low-resolution datasets, all protected models suffer from severe performance degradation with just one error bit. Although the accuracies do not drop to the level of random prediction, they are unacceptably low. When the number of flipped bits has increased to 3, the prediction accuracies of all protected models have dropped to no more than 21%. On the other hand, for ImageNet-10, the accuracies of the models degraded slower with the number of flipped bits. The accuracies dropped by only around 4% with just one bit flip in the user key. However, all four models experienced a performance degradation of more than 40% when 13 key bits are flipped. Considering the image resolution (1024 bits), it is still computationally intractable to successfully brute force 13 correct key bits.
V-B4 Verification and tracing performance
TABLE IV: TSRs (%) obtained by comparing correct keys and decoded keys from intercepted images with the private SteganoGAN decoder.
@@ V-B4 Verification and tracing performance
For each dataset, we select 1,000 benign test images and encode them with 10 different keys, $k_{1},k_{2},\cdots,k_{10}$, to simulate an intercepted dataset $D_{\text{int}}$ that contains 10000 images embedded with unknown keys. These images are fed into the private SteganoGAN decoder to extract their embedded keys. We set the threshold $\epsilon_{3}$ of Eq. (27) to 1, which is the most stringent threshold with at most one bit difference allowed. Table IV demonstrates that an extremely high TSR of over 97% is achievable across all datasets and keys, with an average TSR of above 99% for all five datasets. With such a high TSR, the owner can easily pinpoint the culprit with high confidence, even if only one encoded image is accessible.
TABLE V: Tracing accuracy (%) of the black-box query on 20 suspected models. A suspected model is successfully traced if: 1) it is actually a benign model and identified as innocent; 2) it is correctly identified as a pirated model, and the dishonest buyer is correctly traced.
  Network Dataset
CIFAR-10 CIFAR-100 Caltech-101 GTSRB ImageNet-10
  Res-18 100 100 100 100 100
PreRes-18 100 100 100 100 100
SENet 100 100 100 100 100
Mob-V2 100 100 100 100 100
To evaluate the verification and tracing performance of black-box query, we create a query dataset $D_{q}$ of 100 benign test images and encode them with 10 different keys to produce 10 encoded query datasets, $D^{0}_{q}$, $D^{1}_{q}$, $\cdots$, $D^{9}_{q}$. We also train 10 benign models from scratch using the benign training dataset and mix them with the 10 protected models to create a suspected model set of 10 innocent models and 10 protected models pirated from different users. It is observed that innocent models consistently achieve high performance on all query datasets, regardless of the embedded keys. In contrast, pirated models perform optimally (with 70% to 90% accuracy depending on the dataset) only on a single query dataset embedded with the correct key, and poorly (with $<$ 20% accuracy) otherwise. The tracing accuracy (TA) of black-box query can be computed as:
TA=number of successfully traced modelsnumber of suspected $\text{TA}=\frac{\text{number of successfully traced models}}{\text{number of suspected models}}\times 100\%,$ (28)
Table V shows that 100% TA is achieved across all datasets and networks, which confirms the excellent ownership verification and culprit tracing accuracy of IDEA.

### v2

@@ 8       $\overline{D}_{\text{tr},j}=D_{\text{tr}}+D^{+}_{\text{tr},j}+D^{-}_{\text{tr},j}$
10       Fine-tune the real expert model for $E_{1}$ epochs from the original unprotected model.
11       $\triangleright$ Eq. (11)
@@
13      Separately optimize two randomly initialized models for $E_{2}$ epochs to obtain the fake experts.
14       $\triangleright$ Eq. (12) (13)
15
16      Construct the MoE model $\mathcal{M}$.
17
@@
19       Distill the MoE teacher model for $E_{3}$ epochs to generate the protected student model $f^{\prime}$.
20       $\triangleright$ Eq. (14)
21
22      return $f^{\prime}$.
Distribute each model-key pair and the encoder (i.e., $\{f^{\prime},k\}$ and $\mathcal{E}$) to the corresponding legal user.
@@
IV-C Model Obfuscation
It is worth noting that the MoE model $\mathcal{M}$ lacks a gating network to decide which expert should be activated for prediction. Consequently, $\mathcal{M}$ requires prior knowledge about the domain from which the input image is sampled to activate the correct domain expert, which is impractical in real-world scenarios. Additionally, distributing $\mathcal{M}$ directly to a user poses a security risk, as the adversary can easily identify the real expert due to its superior classification performance on benign test samples compared to the two fake experts. To address these issues, we obfuscate the three experts by unifying them into a single coherent model using knowledge distillation. We use $\mathcal{M}$ to train a student model $f^{\prime}$ which imitates the teacher’s behaviors on not only the fully-connected layer but all selected hidden layers from shallow to deep. To achieve this, $f^{\prime}$ is jointly optimized with three knowledge distillation losses $\mathcal{L}_{\text{kl}}$, $\mathcal{L}_{\text{at}}$ and $\mathcal{L}_{\text{crd}}$ such that
$\mathop{\min}\left[\mathcal{L}_{\text{kl}}+\lambda_{2}\mathcal{L}_{\text{at}}+\lambda_{3}\mathcal{L}_{\text{crd}}\right],$ (14)
where $\lambda_{2}$ and $\lambda_{3}$ are two positive factors to adjust the scale of $\mathcal{L}_{\text{at}}$ and $\mathcal{L}_{\text{crd}}$, respectively.
$\mathcal{L}_{\text{kl}}=\mathbb{E}_{x\sim\mathcal{W}}\left[\mathcal{L}_{\text{KL}}\left(f^{\prime}(x),\mathcal{M}(x)\right)\right]$ (15)
(2) A 3D feature map $f_{l}\in\mathbb{R}^{{c}\times{h}\times{w}}$ can be flattened along the channel dimension as $\mathcal{A}^{\alpha}(f_{l})=\sum^{c}_{i=1}\left|f_{l,(i)}\right|^{\alpha}$, where $\alpha>1$ is the factor that amplifies large neuron activations. Enlarging $\alpha$ places greater emphasis on regions with high output values. $\alpha$ is fixed to 2 according to [47].
The attention transfer loss [48] is computed as:
$\mathcal{L}_{\text{AT}}(f_{l}^{\text{T}},f_{l}^{\text{S}})=\left\|\frac{\mathcal{A}^{\alpha}\left(f_{l}^{\text{T}}\right)}{\left\|\mathcal{A}^{\alpha}\left(f_{l}^{\text{T}}\right)\right\|_{2}}-\frac{\mathcal{A}^{\alpha}\left(f_{l}^{\text{S}}\right)}{\left\|\mathcal{A}^{\alpha}\left(f_{l}^{\text{S}}\right)\right\|_{2}}\right\|_{2},$ (16)
where $f_{l}^{\text{T}}$ and $f_{l}^{\text{S}}$ represent the features extracted from the teacher and student models, respectively at the same layer $l$. We also align the neuron attention of teacher $\mathcal{M}$ and the student $f^{\prime}$ on the selected layers by minimizing
$\mathcal{L}_{\text{at}}=\mathbb{E}_{x\sim\mathcal{W}}\left[\sum_{l\in T_{\text{sel}}}\mathcal{L}_{\text{AT}}\left(\mathcal{M}_{l}(x),f^{\prime}_{l}(x)\right)\right].$ (17)
(3) To ensure that input samples from different domains also share low MI in the latent space in the student model, contrastive representation distillation (CRD) [49] loss is incorporated. Given a tuple of teacher and student representations $(f_{l}^{T}(x_{a}),f_{l}^{S}(x_{b}))$, where $x_{a}$ and $x_{b}$ are two random images sampled from $\mathcal{W}$. It is a positive pair if the two inputs are from the same domain (e.g., $x_{a}\sim\mathcal{A}$ and $x_{b}\sim\mathcal{A}$). Conversely, it is a negative pair if the two inputs are sourced from different domains (e.g., $x_{a}\sim\mathcal{A}$ and $x_{b}\sim\mathcal{B}$). Let $q_{1}$ and $q_{0}$ denote the joint distributions for positive and negative pairs, respectively. The CRD loss is given by:
$\mathcal{L}_{\text{crd}}=$ $-\sum_{l\in T_{\text{sel}}}\mathbb{E}_{q_{1}}\left[\text{log}\left(h_{l}(f_{l}^{T},f_{l}^{S})\right)\right]$ (18)
$-N_{\text{neg}}\sum_{l\in T_{\text{sel}}}\mathbb{E}_{q_{0}}\left[\text{log}\left(1-h_{l}(f_{l}^{T},f_{l}^{S})\right)\right],$ (19)
where $h_{l}(\cdot,\cdot)$ denotes a critic to determine if the two representations are positively paired, and whose parameters can be optimized in parallel with $\mathcal{L}_{\text{crd}}$. We provide one positive pair for every $N_{\text{neg}}$ negative pairs during training, with the aim to pull the positive pairs closer and push the negative pairs apart.
The protected model $f^{\prime}$ has similar performance as the MoE model $\mathcal{M}$, but the functionalities and hidden-layer behaviors of the three experts have been indistinguishably blent into a coherent network, making it infeasible for the adversary to extract the real expert. The algorithm for generating $M$ model-key pairs is depicted in Algorithm 1.
IV-D Verification
Licensee Verification. To securely demonstrate a legitimate licensee without revealing sensitive information, an authorized user provides one or more encoded images instead of directly sharing the secret key. The SteganoGAN decoder, which is kept private by the owner, is used to extract the embedded key from these images. The licensee is validated if the Hamming distance (HD) between the extracted key and a recorded key in the owner’s database falls below a predefined threshold.
Let $D_{\text{e}}$ be an encoded dataset where each image is encoded with a random key, the decoding accuracy (DA) is defined as:
$\text{DA}=\mathbb{E}_{x\sim D_{\text{e}}}\left[\mathbb{I}\{\text{HD}\left(k,\hat{k}\right)\leq\epsilon_{3}\}\right],$ (20)
where $\epsilon_{3}$ is the predefined threshold, $k$ and $\hat{k}$ represent the actual and extracted keys, respectively. A high DA indicates that the decoder can accurately extract embedded keys from encoded images to provide an accurate and secure licensee verification.
Ownership Verification. The owner can conduct a series of black-box queries and analyze the outputs to determine whether $\hat{f}$ is an illegally redeployed model. If unauthorized redistribution is confirmed, the dishonest user responsible can be identified.
Definition 2 (Black-box Query-based Ownership Verification). Let $D^{\mathcal{B}}_{q}$ denote a benign query set sampled from the benign domain. Ownership over the suspected model $\hat{f}$ is verified if and only if there exists exactly one $k\in\mathcal{K}$ satisfying:
$\left|\operatorname{Acc}\{f,D^{\mathcal{B}}_{q}\}-\operatorname{Acc}\{\hat{f},D^{\mathcal{A}}_{q}\}\right|<\epsilon_{1},$ (21)
$\operatorname{Acc}\{f,D^{\mathcal{B}}_{q}\}-\operatorname{Acc}\{\hat{f},D^{\mathcal{N}}_{q}\}>\epsilon_{2},$ (22)
where $D^{\mathcal{A}}_{q}$ consists of query samples embedded with $k$, and $D^{\mathcal{N}}_{q}$ consists of query samples embedded with other recorded keys.
Eq. (21) proves that the full functionality of the suspected model is activated by a specific recorded key, and Eq. (22) confirms that $\hat{f}$ cannot be an innocent model due to its degraded performance on non-matching keys. Ownership is thereby validated, and the dishonest user is identified, as the traced key encodes the identities of both the owner and the user.
TABLE I: Specifications of the experimental datasets.
Dataset # Labels Input Size # Training / Test Images
CIFAR-10 10 32 $\times$ 32 50000 / 10000
GTSRB 43 32 $\times$ 32 39208 / 12630
Caltech-101 101 32 $\times$ 32 6942 / 1735
ImageNet-10 10 224 $\times$ 224 12000 / 1000
@@ V-A1 Datasets and Networks
Table I specifies the four image classification datasets used in our experiments: CIFAR-10 [50], Caltech-101 [51], GTSRB [52], and ImageNet [53], with different resolutions and numbers of classes. To speed up the training without loss of generality, ImageNet evaluation is performed on a subset of 10 randomly selected classes, referred to as ImageNet-10. Four popular architectures are employed in our evaluation: ResNet-18 (Res-18) [54], PreActResNet-18 (PreRes-18) [54], SENet [55] and MobileNet-V2 (Mob-V2) [56].
V-A2 Other settings
The key size parameter $r$ is set to 32 for ImageNet-10, and 16 for the other four datasets. For each classification task, a dedicated SteganoGAN model is trained using 2500 samples randomly selected from the benign training dataset. The unprotected model $f$ is trained from scratch for 200 epochs, and from which the real expert $f_{\text{real}}$ is fine-tuned for 40 epochs, both using the SGD optimizer. Each fake expert is separately trained for 1000 iterations. During the knowledge distillation process, the student model is optimized from scratch for 50 epochs using the Adam optimizer. Ten protected models are generated from the same source model with different random keys for distribution to ten different users. Several common data augmentations, including random cropping, random horizontal flipping, random rotation and random erasing, are applied during training to alleviate overfitting.
@@ V-A2 Other settings
V-B Evaluation
V-B1 Stealth
Fig. 4 presents some examples of benign and authorized images for CIFAR-10, Caltech-101, GTSRB, and ImageNet-10 datasets. We also provide the magnified (×$\times 10$) residual images obtained by the absolute difference in pixel values between the benign and authorized images. The perturbations introduced by SteganoGAN encoding are visually imperceivable. In other words, the attacker can hardly figure out the embedded key by inspecting a few test images encoded with the same key. Furthermore, the perturbations are sample-specific. Even if the attacker can compute a residual image from a pair of benign and authorized images of a protected model, it is impossible to generate an authorized image by adding the residual image to a different test image. The means and standard deviations of three objective image quality assessments (IQA), namely Structural Similarity Index Measure (SSIM) [57], Peak Signal to Noise Ratio (PSNR) and Learned Perceptual Image Patch Similarity (LPIPS) [58], are presented in Table II. The outstanding IQA scores prove the stealth of the key-encoded images.
@@ V-B2 Effectiveness and fidelity
The inference performances of both unprotected and protected models on benign, authorized, and noise test images are presented in Table III. The results indicate that unprotected models consistently achieve high prediction accuracy across all three domains. All protected models fulfill the fidelity requirement of Eq. (4), with inference accuracy of authorized images comparable to the benign accuracy of the corresponding unprotected models. The accuracy drop is no more than 1.5%. Additionally, the protected models perform poorly on benign and noise images, with near-random guessing accuracies, making them useless to unauthorized users without the correct key.
(a) CIFAR-10
(b) Caltech-101
(c) GTSRB
(d) ImageNet-10
Figure 5: Confusion matrices for the evaluation of 10 different keys on the 10 encoded models. The model is ResNet-18.
(a) CIFAR-10
(b) Caltech-101
(c) GTSRB
(d) ImageNet-10
Figure 6: Accuracies of the protected models on test images embedded with incorrect keys with a few erroneous bits.
@@ V-B3 Uniqueness
Ten protected models, each encoded with unique keys, are trained and evaluated across all possible model-key pair combinations. The confusion matrices are presented in Fig. 5, where each row represents a protected model and each column represents a key. The diagonal and off-diagonal values represent test accuracies for matched and mismatched model-key pairs, respectively. The off-diagonal values are consistently low (<20%), indicating that user keys among protected models are non-exchangeable. This ensures precise ownership verification and culprit tracing as a protected model associated with one legitimate user cannot achieve accurate predictions on images encoded with another user’s key.
An attacker may steal a protected model and the SteganoGAN encoder from a legitimate user, but lacks access to the correct key. To evaluate the tolerance of the protected model to bit errors in the key used for encoding the test images, we randomly flipped a few bits, ranging from 1 to 24, of the user keys. As shown in Fig. 6, for the four low-resolution datasets, all protected models suffer from severe performance degradation with just one error bit. Although the accuracies do not drop to the level of random prediction, they are unacceptably low. When the number of flipped bits has increased to 3, the prediction accuracies of all protected models have dropped to no more than 21%. On the other hand, for ImageNet-10, the accuracies of the models degraded slower with the number of flipped bits. The accuracies dropped by only around 4% with just one bit flip in the user key. However, all four models experienced a performance degradation of more than 40% when 13 key bits are flipped. Considering the 1024-bit key length, successfully brute-forcing remains computationally intractable.
TABLE IV: DAs (%) evaluated on an encoded dataset that contains 10,000 images embedded with unknown keys.
@@
V-B4 Licensee Verification
For each dataset, we select 1,000 benign test images and encode them with 10 different keys to simulate an encoded dataset that contains 10,000 images embedded with unknown keys. These images are fed into the owner’s private SteganoGAN decoder to extract their embedded keys. We set the threshold $\epsilon_{3}$ of Eq. (20) to 1, which is an extremely stringent threshold that allows only one bit error. Table IV demonstrates that a high DA of over 97% is achievable across all datasets and keys, with an average DA of above 99% for all five datasets. This level of accuracy enables the owner or a trusted third party engaged by the owner to verify licensees with high confidence, even when provided with just a single encoded image by the user.